OpenStudy (kaedalozzano10):

Which fractions are equivalent to the following? 4/6 and 7/9 A. 4/54 and 7/54 B. 36/54 and 42/54 C. 8/24 and 14/4

OpenStudy (littlequeenasiyah):

c:I think maybe try this http://www.calculatorsoup.com/calculators/math/fractions.php

OpenStudy (faiithhh):

I'd choose b because if you multiply the first fraction by 9, you get 36/54, and if you multiply the second fraction by 6, you get 42/54

OpenStudy (kaedalozzano10):

ok thanks

OpenStudy (welshfella):

You mean 9/9 and 6/6 @faithhh P

OpenStudy (kaedalozzano10):

thank s

OpenStudy (kaedalozzano10):

so the right answer is B

OpenStudy (welshfella):

Yes You have to multiply top and bottom by the same number to check if they are the same.

OpenStudy (welshfella):

4/6 * 9/9 = 36/54

OpenStudy (welshfella):

What you are in effect doing is multiplying by 1 so the new fraction has the same value

OpenStudy (welshfella):

4/6 * 9/9 = 36/54